OpenStudy (waheguru):

i need help fast i have no clue how to solve for this x/6 - 5= 1/2 x

OpenStudy (anonymous):

x-30=3x

OpenStudy (anonymous):

x=-15

OpenStudy (waheguru):

HOW DO I GET THE ANSWER

OpenStudy (anonymous):

multiply everything by 6

OpenStudy (anonymous):

(x/6) = (1/2)x + 5 (x/6) = (3/6)x + 5 -(2/6)x = 5 x = -15

OpenStudy (waheguru):

WHY WHY DO I MILTIPLY BY 6

OpenStudy (anonymous):

to get rid of the fraction

OpenStudy (waheguru):

CAN U SHOW ALL STEPS

OpenStudy (anonymous):

6((1/6)x-5)=6(1/2)x (6/6)x-5*6=(6/2)x x-30=3x x-x-30=3x-x -30=2x -30/2=(2/2)x -15=x

OpenStudy (waheguru):

THX BRO

OpenStudy (anonymous):

no prob